C(s) + 2H2(g) ---> CH4(g)

if the reaction produces 145 liters of CH4, how many liters of H2 were consumed?
1
Expert's answer
2019-03-14T04:01:04-0400

The balanced equation for considered reaction can be written as follows

"C(s) + 2H_{2}(g) \rarr CH_{4}(g)"


"\frac{V(H_{2})}{N(H_{2}) \cdot V_m} = \frac{V(CH_{4})}{N(CH_{4}) \cdot V_m}"

V - volume

N - coefficient in chemical equation

Vm - molar volume

"V(H_{2}) = \frac{V(CH_{4}) \cdot N(H_{2})}{N(CH_{4})}"

"V(H_{2}) = \frac{145 \cdot 2}{1}=290 \; L"


Answer: 290 liters.